OpenStudy (anonymous):

sqrt of x + 2 minus 7 = sqrt of x + 9

OpenStudy (jhannybean):

\[\large \sqrt{x+2}-7 = \sqrt{x+9}\]?

OpenStudy (anonymous):

Yep

OpenStudy (jack1):

first, do no harm 2nd, square everything what'd jew get?

OpenStudy (anonymous):

Do I square the 7 too?

OpenStudy (jhannybean):

\[\large \sqrt{x+2}-7 = \sqrt{x+9}\] square everything on both sides to get rid of the square root \[\large (\sqrt{x+2}-7)^2=(\sqrt{x+9})^2\]\\[\large (\sqrt{x+2}-7)(\sqrt{x+2}-7)=x+9\]foil method \[\large (\sqrt{x+2})(\sqrt{x+2})-(\sqrt{x+2})(7)-(7)(\sqrt{x+2})-(7)(-7)=x+9\] Simplify \[\large x+2-7\sqrt{x+2}-7\sqrt{x+2}+49=x+9\]

OpenStudy (jack1):

@Jhannybean that seems kinda longhand ... why not:

OpenStudy (jhannybean):

haha that's how i am able to check D:

OpenStudy (jhannybean):

after im done :(

OpenStudy (anonymous):

lol

OpenStudy (jhannybean):

\[(\sqrt{x+2})(\sqrt{x+2})-(\sqrt{x+2})(7)-(7)(\sqrt{x+2})-(7)(-7)=x+9\]\[x+2-14\sqrt{x+2}+49=x+9\]\[x-14\sqrt{x+2}+52=x+9\]\[-14\sqrt{x+2}=-43\] bleh.

OpenStudy (jhannybean):

Shall i just....delete my posts...<_<

OpenStudy (anonymous):

no!!

OpenStudy (anonymous):

I'm definitely copying as we speak, lmao

sam (.sam.):

No solution

OpenStudy (jhannybean):

oh LOL

OpenStudy (anonymous):

How is it no solution O_o

OpenStudy (anonymous):

Ohmahgerd D;

OpenStudy (jhannybean):

@.Sam. just throws out an answer :P

OpenStudy (anonymous):

Yes, so helpful. Not.

OpenStudy (jack1):

yeah, my math didn't work out... thought i'd done something wrong hey sorry @Jhannybean ... Jack's bad

OpenStudy (anonymous):

Hey @Jhannybean what would the final answer be for x or did you get it lol

OpenStudy (jhannybean):

lol what? xD i think @Jack1 has it :P

OpenStudy (jack1):

no it doesnt work LHS = x + 51 - (14(sqrt(x+2)) RHS = x+9

OpenStudy (jhannybean):

continuing from \[\large -14\sqrt{x+2}=-43\]\[\large \sqrt{x+2}=\frac{43}{14}\] square both sides \[(\sqrt{x+2})^2=(\frac{43}{14})^2\]\[\large x+2=\frac{1849}{196} \implies x= \frac{1457}{196}\] that's not right, i believe.

OpenStudy (jhannybean):

I don't even know anymore.

OpenStudy (jack1):

but does that work again in the initial eqn?

OpenStudy (jack1):

sqrt of x + 2 minus 7 = sqrt of x + 9 -55/14 = 4.0538 ish -3.929 does not equal 4.0538 ish

OpenStudy (jhannybean):

OMG SO CLOSE.STUPID NEGATIVE SIGN.

sam (.sam.):

I don't think that's the answer, try substituting it in, see if LHS = RHS

OpenStudy (jack1):

... i just did that

OpenStudy (jack1):

why do u hate me @.Sam. ...?

OpenStudy (jhannybean):

xDDD

sam (.sam.):

I didn't, I just have a different thinking :)

OpenStudy (jack1):

is it coz i'm black?

OpenStudy (jack1):

:)

sam (.sam.):

Lol, I don't even know that

sam (.sam.):

and what's wrong with black?

OpenStudy (jhannybean):

LOL REVELATION?

OpenStudy (anonymous):

Lmao okay next problem is sqrt of 2x+1 + qrt of 2x+6 = 5

sam (.sam.):

\[[(\sqrt{x+2}-\sqrt{x+9})]^2=7^2 \\ \\ (x+2)-2\sqrt{(x+2)(x+9)}+(x+9)=49 \\ \\ 2x+11-2\sqrt{(x+2)(x+9)}=49 \\ \\ 2(x-\sqrt{(x+2)(x+9)})+11=49 \\ \\ x-\sqrt{(x+2)(x+9)}=19 \\ \\ (x-19)^2=(x+2)(x+9) \\ \\ 20=58x \\ \\ x=\frac{10}{29}\]

sam (.sam.):

See, \(LHS \neq RHS\)

OpenStudy (saifoo.khan):

Jaan needs a sleep. She has been making mistakes all day.

OpenStudy (jhannybean):

=_=

OpenStudy (jack1):

she just wrote like 3 pages of squiggly equation posts, she deserves shiny shiny medals

OpenStudy (jack1):

@randomguy111 prettys sure your post's been hijacked, may want to close and pop ur new q in a new post...?